# Low-resolution Prior Equilibrium Network for CT Reconstruction

### v1

@@ Abstract
The unrolling method has been investigated for learning variational models in X-ray computed tomography. However, it has been observed that directly unrolling the regularization model through gradient descent does not produce satisfactory results. In this paper, we present a novel deep learning-based CT reconstruction model, where the low-resolution image is introduced to obtain an effective regularization term for improving the network‘s robustness. Our approach involves constructing the backbone network architecture by algorithm unrolling that is realized using the deep equilibrium architecture. We theoretically discuss the convergence of the proposed low-resolution prior equilibrium model and provide the conditions to guarantee convergence. Experimental results on both sparse-view and limited-angle reconstruction problems are provided, demonstrating that our end-to-end low-resolution prior equilibrium model outperforms other state-of-the-art methods in terms of noise reduction, contrast-to-noise ratio, and preservation of edge details.
Keywords: CT reconstruction, limited-angle, sparse-view, deep equilibrium model, prior-based regularization, low-resolution image
@@ 1 Introduction
Computed Tomography (CT) is a fundamental imaging tool that finds wide applications in various fields, such as industrial non-destructive testing, medical diagnoses, and security inspections. Specifically, let us consider the observation of a corrupted set of measurements denoted as $\bm{b}$, obtained by applying a linear measurement operator $\bm{A}$ to an image $\bm{x}$ and introducing certain additive noises $\epsilon$. The relationship can be expressed as follows
$\bm{A}\bm{x}+\epsilon=\bm{b}.$ (1)
@@
Our objective is to reconstruct the clean image $\bm{x}$ from the provided measurements $\bm{b}$ and the system matrix $\bm{A}$. In general, when the number of pixels in a reconstructed image exceeds the number of projection samples in CT imaging, the inverse problem of (1) becomes ill-posed.
@@
Due to the radiation, various approaches have been explored, which can be primarily categorized into two main categories, either to modify the scanning protocol by reducing the tube voltage and/or tube current [1] or downsample the measured data for CT reconstruction, such as interior CT [2, 3, 4, 5], and sparse-view CT [6, 7, 8, 9, 10, 11]. These approaches aim to achieve dose reduction while maintaining satisfactory image quality. However, employing the first category of approaches for radiation dose reduction may result in the measured CT data being heavily contaminated with noise. The standard filtered back-projection (FBP) method, when applied to reconstruct CT images using noise-contaminated CT data, suffers from a significant degradation in image quality. On the other hand, the second category of approaches, aimed at reducing radiation dose, usually introduces additional ill-posedness into the inverse problem. It is worth noting that the regularization method is an effective technique for tackling ill-posed problems by incorporating prior information. The total variation (TV) regularization, as a prominent example, is considered state-of-the-art in handling low-dose (LD) and few-view CT [12]. Moreover, alternative sparsity regularization techniques, e.g., the utilization of wavelet frames [13], have been investigated in CT reconstruction. A noise suppression-guided image filtering reconstruction algorithm has recently been proposed as a solution to address the low signal-to-noise ratio (SNR) problem [14]. With recent advancements in artificial intelligence and hardware performance, deep learning (DL) has emerged as a promising approach for denoising in low-dose computed tomography, which has demonstrated encouraging results. Since then, various deep learning reconstruction methods have been studied including the pre-processing in the projection domain, projection-to-image reconstruction, and post-processing in the image domain [15, 16, 17, 18, 19, 20, 21]. These studies have shown that DL approaches consistently deliver enhanced or comparable noise suppression and maintain structural fidelity [22, 23, 24, 25, 26].
1.1 Prior works
One remaining issue is the theoretical analysis of the learning-based reconstruction methods, especially the convergence analysis. Broadly speaking, we can roughly divide the learning-based methods into three categories as the post-processing methods [27, 28], the Plug-and-play priors methods [29, 30, 31, 32, 33] and the unrolling methods [34, 35].
@@ 1.1 Prior works
Post-processing methods. A deep neural network-based post-processing of a model-based reconstruction is parameterized as $\mathcal{R}_{\theta}:=\mathcal{P}_{\theta}\circ\mathcal{B}$, where $\mathcal{B}$ denotes a classical reconstruction method (e.g., FBP) and $\mathcal{P}_{\theta}$ represents a deep convolutional network with parameters $\bm{\theta}$. In general, there is no convergence guarantee for the post-processing strategies. Indeed, a small value of $\|\bm{A}\bm{x}-\bm{b}\|$ does not necessarily imply a small value of $\|\bm{A}\mathcal{P}_{\theta}(\bm{x})-\bm{b}\|$ for the output of $\mathcal{P}_{\theta}$. Such issue was addressed in [27] by parametrizing the operator $\mathcal{P}_{\theta}$ as $\mathcal{P}_{\theta}=\bm{I}+\left(\bm{I}-\bm{A}^{\top}\bm{A}\right)\mathcal{Q}_{\theta}$, where $\mathcal{Q}_{\theta}$ is a Lipschitz-continuous DNN. Since $\bm{I}-\bm{A}^{\top}\bm{A}$ is the projection operator in the null-space of $\bm{A}$, the operator $\mathcal{P}_{\theta}$ always satisfies $\bm{A}\mathcal{P}_{\theta}\left(\bm{x}\right)=\bm{A}\bm{x}$. Null-space networks are demonstrated to offer convergent regularization schemes. Inspired by the theory of optimal transport, an adversarial framework for learning the regularization term was proposed in [28], which possesses the stability guarantee, subjecting to the condition that the regularization is 1-Lipschitz and coercive.
@@
Plug-and-play (PnP) priors methods. The PnP prior approach leverages the strengths of both - realistic modeling through physical knowledge and flexible learning from data patterns. Relying on the maximum a posteriori probability (MAP) estimator, the solution of (1) can be formulated as the following minimization problem
${{\bm{x}}}\in\mathop{\arg\min}\limits_{\bm{x}}\mathcal{S}(\bm{x},\bm{b})+\mathcal{R}(\bm{x};\bm{\theta}),$ (2)
where $\mathcal{S}(\bm{x},\bm{b})=-\log(p_{{b}|{x}}(\bm{x}))$ is the likelihood relating the solution $\bm{x}$ to the measurement $\bm{b}$ and $\mathcal{R}(\bm{x};\bm{\theta})=-\log(p_{x}(\bm{x}))$ is the prior distribution. The PnP priors framework models the likelihood by obeying the noise distributions of the underlying physical process, while it extracts flexible priors from data observation. The seminal study conducted by Venkatakrishnan et al. [36] introduced the innovative concept of substituting denoising algorithms as priors for model-driven inversion, where the denoising algorithm was used inside the Projected Gradient Descent to obtain an iterative scheme. Here, the choice of denoiser can follow model-based approaches such as BM3D, but in recent years, deep neural network-based denoisers have been increasingly widely used in various inverse problems. These methods have shown excellent empirical performance, which inspired the theoretical studies on its convergence.
@@
The first result demonstrating the global objective convergence of PnP-ADMM was presented in [29], where the denoiser used in the PnP-ADMM algorithm should satisfy two requirements: it must be continuously differentiable, and its gradient matrix should be doubly stochastic. The fixed-point convergence of PnP-ADMM with a continuation scheme and bounded denoisers was established in [30]. Ryu et al. [31] demonstrated that the iterations of both PnP-proximal gradient-descent and PnP-ADMM exhibit contraction behavior when the denoiser satisfies the Lipschitz condition. The fixed-point convergence of PnP-forward-backward splitting (FBS) and PnP-ADMM were proven in [32]. Specifically, the proof was established for linear denoisers of the form $Denoiser(\bm{x})=\bm{Wx}$, where $\bm{W}$ is diagonalizable and its eigenvalues are within the range of [0, 1]. Convergence guarantees for PnP methods were derived in [37] with gradient-step (GS) denoisers, alleviating the need for such restrictive assumptions. The GS denoisers are constructed as $Denoiser=\bm{I}-\nabla g_{\sigma}$, where $g_{\sigma}(x)=\|\bm{x}-{P}_{\sigma}(\bm{x})\|$, with ${P}_{\sigma}$ being a deep network without any structural constraints. This parametrization was shown to have enough expressive power to achieve state-of-the-art denoising performance in [37]. In [33], a closely related PnP approach was employed with a similar objective of providing an asymptotic characterization of the iterative PnP solutions. The key concept was to model maximally monotone operators (MMO) using a neural network, where the parameterization of MMOs was achieved by modeling the resolvent through a non-expansive neural network.
@@
$\eqalign{\mathop{\min}\limits_{\bm{\theta}}&~{}~{}\frac{1}{m}\sum_{i=1}^{m}\ell\left({x}_{i}(\bm{\theta}),{x}^{*}_{i}\right)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\hfill(\textup{UL})\\ \mbox{s.t.}&~{}~{}{x}_{i}(\bm{\theta})\in\mathop{\arg\min}\limits_{\bm{x}}\mathcal{S}\left(\bm{x},{b}_{i}\right)+\mathcal{R}(\bm{x};\bm{\theta})\quad\quad\quad\quad\quad\hfill(\textup{LL})}$ (3)
@@
where a loss function $\ell$ is employed to enforce the similarity between the reconstructed image $\bm{x}$ and ground truth ${\bm{x}}^{*}$. Figure 1 illustrates a generic bilevel problem for image reconstruction. The upper-level (UL) loss function evaluates the quality of a vector of learnable parameters, while it also depends on the solution to the lower-level (LL) cost function.
@@
The bilevel problems [40, 41, 42, 43, 44] are known for their inherent difficulty to be solved numerically. To utilize gradient descent methods, it is necessary to compute the derivative of the solution operator for the lower-level problem (2) with respect to the parameters $\bm{\theta}$. If the objective of the lower-level problem is differentiable and has a unique minimum, gradients can be computed using implicit differentiation [45]. However, non-smooth objectives, along with potentially multi-valued solution operators, are frequently employed in image and signal processing. The method of unrolling involves utilizing an iterative algorithm that solves the lower-level problem and replaces the optimal solution ${{x}}_{i}$ with the $N$-th iteration. Let us denote the iteration by $F_{\theta}^{(N-1)}(\cdot,\bm{b};\bm{\theta})$. The lower-level problem by unrolling amounts to solve the surrogate problem as follows
$\min_{\bm{\theta}}~{}~{}\frac{1}{m}\sum_{i=1}^{m}\ell\left(F_{\theta}^{(N-1)}\left(x_{i},{b}_{i};\bm{\theta}\right),{x}^{*}_{i}\right).$
Subsequently, the same algorithmic scheme $F_{\theta}^{(N-1)}(\cdot,\bm{b};\bm{\theta})$ with the learned parameters ${{\bm{\theta}}}$ to make predictions on unseen data. We can express the unrolling network as follows
$\bm{x}^{(k+1)}=F_{\theta}^{(k)}\left(\bm{x}^{(k)},\bm{b};\bm{\theta}\right)\quad\mbox{for}\quad k=0,\ldots,N-1,$
@@
where $k$ is the iteration index and $F_{\theta}^{(k)}(\cdot,\bm{b};\bm{\theta})$ represents a nonlinear transformation, such as inner products followed by the application of a nonlinear activation function. In fact, different iterations can share the same weights, such that $F_{\theta}^{(k)}(\cdot,\bm{b};\bm{\theta})$ at each iteration remains the same, i.e., $F_{\theta}^{(k)}(\cdot,\bm{b};\bm{\theta})=F_{\theta}(\cdot,\bm{b};\bm{\theta})$ for all $k$. The equilibrium or fixed-point-based networks introduced in [34, 35] introduced weight-sharing into the unrolling methods, which has demonstrated the capability to attain competitive performance. The main concept was to represent the output of a feed-forward model as a fixed-point of a nonlinear transformation, which enables the use of implicit differentiation for back-propagation. Let us consider an $N$-iteration network with input $\bm{b}$ and weights $\bm{\theta}$, and the output of the $k$-th hidden layer to be $\bm{x}^{(k)}$, we have the following recursion
$\bm{x}^{(k+1)}=F_{\theta}\left(\bm{x}^{(k)},\bm{b};\bm{\theta}\right).$ (4)
@@
The limit of $\bm{x}^{(K)}$ as $K\rightarrow\infty$, provided it exists, is a fixed point of the operator $F_{\theta}(\cdot,\bm{b};\bm{\theta})$. This series converges if the Jacobian $\partial_{x}F_{\theta}$ has spectral norm strictly less than 1 when evaluated at any $\bm{x}$, which is true if and only if the iteration map $F_{\theta}$ is contractive. The deep equilibrium model (DEM) serves as a bridge between conventional fixed-point methods in numerical analysis and learning-based techniques for solving inverse problems.

### v2

@@ Abstract
The unrolling method has been investigated for learning variational models in X-ray computed tomography. However, for incomplete data reconstruction, such as sparse-view and limited-angle problems, the unrolling method of gradient descent of the energy minimization problem cannot yield satisfactory results. In this paper, we present an effective CT reconstruction model, where the low-resolution image is introduced as a regularization for incomplete data problems. In what follows, we utilize the deep equilibrium approach to unfolding of the gradient descent algorithm, thereby constructing the backbone network architecture for solving the minimization model. We theoretically discuss the convergence of the proposed low-resolution prior equilibrium model and provide the necessary conditions to guarantee its convergence. Experimental results on both sparse-view and limited-angle reconstruction problems are provided, demonstrating that our end-to-end low-resolution prior equilibrium model outperforms other state-of-the-art methods in terms of noise reduction, contrast-to-noise ratio, and preservation of edge details.
Keywords: CT reconstruction, limited-angle, sparse-view, deep equilibrium model, unrolling, low-resolution image prior
@@ 1 Introduction
Computed Tomography (CT) is a fundamental imaging tool that finds wide applications in various fields, such as industrial non-destructive testing, medical diagnoses, and security inspections. Specifically, let us consider the observation of a corrupted set of measurements, i.e., $\bm{b}\in\mathbb{R}^{N_{views}\times N_{bins}}$, obtained by applying a linear measurement operator $\bm{A}:\mathbb{R}^{N^{2}}\rightarrow\mathbb{R}^{N_{views}\times N_{bins}}$ to a target $\bm{u}\in\mathbb{R}^{N^{2}}$ with certain additive noises $\xi\in\mathbb{R}^{N_{views}\times N_{bins}}$. Mathematically, we can formulate the inverse problem as follows
$\bm{A}\bm{u}+\xi=\bm{b},$ (1)
@@
where $N_{{views}}$ represents the number of angles within the angular interval, $N_{{bins}}$ represents the number of units on the detector, and $N$ represents the denotes the length and width of the target image. Our objective is to reconstruct the clean image $\bm{u}$ from the measurement $\bm{b}$. In general, when the number of pixels in a reconstructed image exceeds the number of projection samples in CT imaging, the inverse problem of (1) becomes ill-posed.
@@
Due to the radiation, various approaches have been explored, which can be primarily categorized into two main categories, either to modify the scanning protocol by reducing the tube voltage and/or tube current [44] or downsample the measured data for CT reconstruction, such as interior CT [36, 51, 60, 70], and sparse-view CT [68, 66, 34, 63, 72, 65]. These approaches aim to achieve dose reduction while maintaining satisfactory image quality. However, employing the first category of approaches for radiation dose reduction may result in the measured CT data being heavily contaminated with noise. The standard filtered back-projection (FBP) method, when applied to reconstruct CT images using noise-contaminated CT data, suffers from a significant degradation in image quality. On the other hand, the second category of approaches, aimed at reducing radiation dose, usually introduces additional ill-posedness into the inverse problem. It is worth noting that the regularization method is an effective technique for tackling ill-posed problems by incorporating prior information. The total variation (TV) regularization, as a prominent example, is considered state-of-the-art in handling low-dose (LD) and few-view CT [38]. Moreover, alternative sparsity regularization techniques, e.g., the utilization of wavelet frames [69], have been investigated in CT reconstruction. A noise suppression-guided image filtering reconstruction algorithm has recently been proposed as a solution to address the low signal-to-noise ratio (SNR) problem [31].
With recent advancements in artificial intelligence and hardware performance, deep learning (DL) has emerged as a promising approach for denoising in low-dose computed tomography, which has demonstrated encouraging results. Since then, various deep learning reconstruction methods have been studied including the pre-processing in the projection domain, projection-to-image reconstruction, and post-processing in the image domain [19, 25, 28, 2, 46, 49, 50]. These studies have shown that DL approaches consistently deliver enhanced or comparable noise suppression and maintain structural fidelity [57, 37, 54, 59, 3]. Broadly speaking, we can roughly divide the learning-based methods for CT reconstruction into three categories as the post-processing methods [56, 40], the plug-and-play priors methods [61, 10, 55, 48, 52] and the unrolling methods [4, 22]. There methods have greatly improved the reconstruction qualities compared to conventional methods. The remaining issues mainly include: firstly, the theoretical property guarantees of learning-based reconstruction methods have not been fully addressed, such as the convergence of the learning-based models. Secondly, for severely ill-posed reconstruction problems, such as limited-angle and sparse-view reconstruction issues, it still requires effective learning-based solutions.
1.1 Convergence studies of learning-based methods
@@
Post-processing methods. A deep neural network-based post-processing of a model-based reconstruction is parameterized as $\mathcal{P}_{\theta}\circ\mathcal{B}$, where $\mathcal{B}$ denotes a classical reconstruction method (e.g., FBP), $\mathcal{P}_{\theta}$ represents a deep convolutional network with parameters $\bm{\theta}$ and $\circ$ represents composite operations. In general, there is no convergence guarantee for the post-processing strategies. That means, a small value of $\|\bm{A}\bm{u}-\bm{b}\|$ does not necessarily imply a small value of $\|\bm{A}\mathcal{P}_{\theta}(\bm{u})-\bm{b}\|$ for the output of $\mathcal{P}_{\theta}$. Such issue was addressed in [56] by parametrizing the operator $\mathcal{P}_{\theta}$ as $\mathcal{P}_{\theta}=\bm{I}+\left(\bm{I}-\bm{A}^{\top}\bm{A}\right)\mathcal{Q}_{\theta}$, where $\mathcal{Q}_{\theta}$ is a Lipschitz-continuous deep convolutional neural network. Since $\bm{I}-\bm{A}^{\top}\bm{A}$ is the projection operator in the null-space of $\bm{A}$, the operator $\mathcal{P}_{\theta}$ always satisfies $\bm{A}\mathcal{P}_{\theta}\left(\bm{u}\right)=\bm{A}\bm{u}$. Null-space networks were demonstrated to offer convergent regularization schemes. Inspired by the theory of optimal transport, an adversarial framework for learning the regularization term was proposed in [40], which possessed the stability guarantee, subjecting to the condition that the regularization is 1-Lipschitz and coercive.
@@
Plug-and-play (PnP) priors methods. The PnP prior approach leverages the strengths of both realistic modeling through physical knowledge and flexible learning from data patterns. Relying on the maximum a posteriori probability (MAP) estimator, the solution of (1) can be formulated as the following minimization problem
${{\bm{u}}}\in\mathop{\arg\min}\limits_{\bm{u}}\mathcal{S}(\bm{u},\bm{b})+\mathcal{R}(\bm{u};\bm{\theta}),$ (2)
@@
where $\mathcal{S}(\bm{u},\bm{b})=-\log(p_{{\bm{b}}|{\bm{u}}}(\bm{u}))$ is the likelihood relating the solution $\bm{u}$ to the measurement $\bm{b}$ and $\mathcal{R}(\bm{u};\bm{\theta})=-\log(p_{\bm{u}}(\bm{u}))$ is the denoiser to be learned by deep neural network models. The PnP priors framework models the likelihood by obeying the noise distributions of the underlying physical process, while it extracts flexible priors from data observation. The learning-based regularization have shown excellent empirical performance, which inspired the theoretical studies on its convergence. The first result demonstrating the global objective convergence of PnP-ADMM was presented in [61], where the learned denoiser $\mathcal{R}(\bm{u};\bm{\theta})$ should satisfy two requirements, i.e., it is continuously differentiable, and its gradient matrix is doubly stochastic. The fixed-point convergence of PnP-ADMM with a continuation scheme and bounded denoiser was established in [10]. Ryu et al. [55] demonstrated that the iterations of both PnP-proximal gradient-descent and PnP-ADMM exhibit contraction behavior when the denoiser $\mathcal{R}(\bm{u};\bm{\theta})$ satisfies the Lipschitz condition. The fixed-point convergence of PnP-forward-backward splitting (FBS) and PnP-ADMM were proven in [48]. Specifically, the proof was established for linear denoisers of the form $\mathcal{R}(\bm{u};\bm{\theta})=\bm{Wu}$, where $\bm{W}$ is diagonalizable and its eigenvalues are within the range of [0, 1]. The convergence guarantees for the PnP methods were derived in [33] with gradient-step (GS), alleviating the need for such restrictive assumptions. The GS denoiser, denoted as $\mathcal{R}(\bm{u};\bm{\theta})=\bm{I}-\nabla g_{\sigma}$, is formulated with $g_{\sigma}(\bm{u})=|\bm{u}-{P}_{\sigma}(\bm{u})|$, where ${P}_{\sigma}$ represents a deep neural network devoid of structural constraints. The parametrization was shown to have enough expressive power to achieve state-of-the-art denoising performance in [33]. In [52], a closely related PnP approach was employed with the objective of providing an asymptotic characterization of the iterative PnP solutions. The key concept was to model maximally monotone operators (MMO) using a deep neural network, where the parameterization of MMOs was achieved by modeling the resolvent through a non-expansive neural network.
@@
$\eqalign{\mathop{\min}\limits_{\bm{\theta}}&~{}~{}\frac{1}{m}\sum_{i=1}^{m}\ell\left(\bm{u}_{i}(\bm{\theta}),\bm{u}^{*}_{i}\right)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\hfill(\textup{UL})\\ \mbox{s.t.}&~{}~{}\bm{u}_{i}(\bm{\theta})\in\mathop{\arg\min}\limits_{\bm{u}}\mathcal{S}\left(\bm{u},{b}_{i}\right)+\mathcal{R}(\bm{u};\bm{\theta})\quad\quad\quad\quad\quad~{}\hfill(\textup{LL})}$ (3)
@@
where a loss function $\ell(\cdot,\cdot)$ is employed to enforce the similarity between the reconstructed image $\bm{u}_{i}$ and ground truth ${\bm{u}_{i}}^{*}$, $\mathcal{S}$ represents the data term, and $\mathcal{R}$ denotes the learnable regularization term, $m$ indicates the number of training data samples. The upper-level (UL) loss function evaluates the quality of a vector of learnable parameters, while it also depends on the solution to the lower-level (LL) cost function. Figure 1 illustrates a generic bilevel problem for image reconstruction, where the model-based reconstruction represents the process of solving the lower-level minimization problem.
@@
The bilevel problems [5, 7, 13, 16, 15] are known for their inherent difficulties to be solved numerically. To utilize gradient descent methods, it is necessary to compute the derivative of the solution operator for the lower-level problem (2) with respect to the parameters $\bm{\theta}$. If the objective of the lower-level problem is differentiable and has a unique minimum, gradients can be computed using implicit differentiation [18]. However, non-smooth objectives, along with potentially multi-valued solution operators, are frequently employed in image and signal processing tasks. The method of unrolling involves utilizing an iterative algorithm that solves the lower-level problem and replaces the optimal solution ${\bm{u}}_{i}(\bm{\theta})$ with the $K$ stages. We can express the unrolling network as follows
$\bm{u}^{(k+1)}=f_{\theta}^{(k)}\left(\bm{u}^{(k)},\bm{b};\bm{\theta}\right)\quad\mbox{for}\quad k=0,\ldots,K-1,$ (4)
where $k$ is the stage index and $f_{\theta}^{(k)}(\cdot,\bm{b};\bm{\theta})$ represents a nonlinear transformation, such as convolution followed by the application of a nonlinear activation function.
@@
In fact, different stages can share the same weights, such that $f_{\theta}^{(k)}(\cdot,\bm{b};\bm{\theta})$ at each stage remains the same, i.e., $f_{\theta}^{(k)}(\cdot,\bm{b};\bm{\theta})=f_{\theta}(\cdot,\bm{b};\bm{\theta})$ for all $k$. The equilibrium or fixed-point-based networks introduced in [4, 22] used weight-sharing into the unrolling methods, which has demonstrated the capability to attain competitive performance. The main concept was to represent the output of a feed-forward neural network model as a fixed-point of a nonlinear transformation, which enabled the use of implicit differentiation for the back-propagation. Let us consider the $K$-stage network model with the input $\bm{b}$ and weights $\bm{\theta}$. The output of the $(k+1)$-th hidden layer can be denoted as $\bm{u}^{(k+1)}$ obtained by the following recursion process
$\bm{u}^{(k+1)}=f_{\theta}\left(\bm{u}^{(k)},\bm{b};\bm{\theta}\right).$ (5)
@@
The limit of $\bm{u}^{(K)}$ as $K\rightarrow\infty$, provided it exists, is a fixed point of the operator $f_{\theta}(\cdot,\bm{b};\bm{\theta})$. The series converges if the spectral norm of the Jacobian $\partial_{u}f_{\theta}$ is strictly less than 1 for any $\bm{u}$, which is true if and only if the iteration map $f_{\theta}$ is contractive. The deep equilibrium model (DEM) serves as a bridge between the conventional fixed-point methods in numerical analysis and learning-based techniques for solving the inverse problems. We illustrate the general diagram of the unrolling algorithm based on gradient descent of the low-level minimization problem in Figure 2, while represent the structure of deep equilibrium model in Figure 3. As can be shown, the difference between deep unrolling model and deep equilibrium model lies in the iterative regime, where deep equilibrium model shares the weights for all stage. Indeed, we use the deep neural networks to learn the gradient operator of the data fidelity and regularization used in the gradient descent scheme.
Figure 2: Depiction of a typical unrolling method using the gradient decent.
Figure 3: Depiction of a typical deep equilibrium method using the gradient decent.